if a 4 +1\a 4 =727,then the value of a 3 -1\a 3 is. How to solve it by using formulae

if a4+1\a4=727,then the value of a3-1\a3 is. How to solve it by using formulae

Grade:9

1 Answers

Vikas TU
7 years ago
a4+1\a4=727
this the given eqn.
Now add 2 both seds so that we can use the suitable identity.
a^4 + 1/a^4 + 2 = 729
(a^2)^2 + 1/(a^2)^2 + 2a^2/a^2 = 27^2
(a^2 + 1/a^2)^2 = 27^2
(a^2 + 1/a^2) = 27
like this now sutract 2 both sides .
R.H.D would become sq. of 5 that is 25.
And u would get,
a – 1/a = 5
now then take cube both sides
and now substitute also (x – 1/x).